# Consumption Smoothing in Metropolis: Evidence from the Working-class Households in Prewar Tokyo

### v8

@@ 6 Conclusion
This study investigates the consumption smoothing behaviors among working-class households in Tokyo. I digitized the daily longitudinal budget survey on the factory worker households conducted in Tsukishima and implemented a systematic econometric method to obtain comparable estimates for income elasticities. The estimated elasticity for total consumption indicates that the factory worker households suffered certain consumption losses when they faced idiosyncratic income shocks. Despite this, I found evidence that while expenses on dispensable categories are likely to be cut in response to the income losses, expenses on indispensable categories are partially smoothed.
The estimated elasticity for total expenditure is in a similar range to those in rural economies in developing countries, although only a few available references exist. According to Ravallion and Chaudhuri (1997), the elasticity for the rural households in Indian villages between 1975 and 1984 was roughly $0.5$. As for the urban households, the estimated elasticity is $0.4$ for Bangkok between 1975 and 1990, which had a similar GDP per capita to 1920 Japan (Townsend 1995). Since my estimate is from a specific occupation, skilled factory workers, one must be careful when making comparisons with these estimates based on urban samples, which include a variety of occupations. At least, however, the factory workers in the historical metropolis of Japan around WWI had been relatively vulnerable to the idiosyncratic income shocks.
A recent study found an income elasticity of $0.39$ for factory workers’ households in Osaka city circa 1920 (Ogasawara 2024). My estimate from the adjusted monthly panels is slightly larger than this estimate. Although the measurement errors in aggregate consumption and income in the monthly panel dataset may explain the discrepancy in these estimates, if taken literally, this implies that the Tsukishima factory workers were more vulnerable than those in Osaka. This indicates that higher levels of urbanization do not necessarily correspond to a greater availability of risk-mitigation institutions. Therefore, inter-city and cross-national comparisons of income elasticities and analysis of the institutional characteristics behind the consumptions there would be another subject of future work.
The finding on the attenuation bias due to measurement errors is potentially important because it may prompt a reinterpretation of existing results. Errors in the timing of payroll can result in inconsistencies in the aggregate data for income and expenditure, albeit the household-level studies often use aggregate data. In this case, modifying the aggregation period to align with wage payments can reduce attenuation bias. Even if the aggregation period is unadjustable, clarifying payroll timing and discussing potential bias is useful.
The results from mechanism analyses provide suggestive evidence that credit purchases played an important role in mitigating adverse idiosyncratic shocks on consumption. Although the institutional details of credit purchases should be investigated further, this finding offers a new historical view of the interdependence between consumers and local retailers. Understanding the local retail system would be important in comprehending regional consumption behavior and, in turn, in understanding the improvement in living standards in prewar Japan.
@@ A.1 Developments of the Cities
Figure A.1: Developments of the Large Cities in Japan
Note: This figure illustrates the development of large cities in Japan between 1889 and 1920. Figure A.1a shows the number of people in Tokyo, Osaka, Kyoto, and Nagoya cities. Figure A.1b shows the share of people in Tokyo (diamond) and the cities with more than $50,000$ people (circle) in total population in Japan.
@@
1. Panel A: The number of people and male workers measured in the Population Census conducted in October 1920 are listed.
2. Panel B: Occupations are classified based on the industrial classification of the 1920 Population Census. Unemployed people are not included in the calculation.
3. Panel C: The number of electrical lights measured at the end of 1921 is divided by the number of people in the 1920 Population Census. The number of houses accessible to water supply measured at the end of September 1920 is divided by the number of total households measured in the 1920 Population Census (%).
4. Panel D: The average daily temperature and precipitation between 1st January 1918 and 31st December 1923 and between 1st February 1920 and 30th April 1920 are listed.
@@ A.2 Financial Institutions in the Cities
Several financial institutions were available for working-class households in the cities around 1920.
@@
A representative savings institution is the postal savings (yūbin chokin). Postal savings was initially designed as a small savings institution for low-income households, and the share of postal savings to total savings was roughly $10$% in the 1910s (Tanaka 2018, p. 29). However, the number of postal savings accounts per total population in the early 1910s exceeded $20$% (Tanaka 2018, p. 33). [^86] James and Suto (2011) show that the saving rate in prewar Japan was higher than that of the U.S. As shown in Panel A of Table A.2, roughly $1.4$ million people had accounts in Tokyo in 1920. This covers approximately $67$% of the entire population in Tokyo and the average savings amount per capita was $64.5$ yen. In Kyobashi Ward, the average figure for the postal savings accounts was $44.5$ yen at the end of 1919 (Tokyo City Office 1921, pp. 892–893). This amount is less than the average monthly earnings of the skilled-factory workers (Section 3.2). Unfortunately, the statistics on postal savings by depositors’ occupation are unavailable. [^87] There are some statistics for the entire Japan. According to the Postal Savings Bureau (1924, p. 72), the share of depositors in the manufacturing sector was $4.8$% (the share based on the deposit amount is $5.5$%). However, this implies that, on average, postal savings may not provide sufficient temporary income.
@@
Savings bank (chochiku ginkō) is another famous savings institution for working-class households in the 1910s (Ito and Saito 2019, pp 79–81). Panel A of Table A.2 shows that the number of people having savings accounts was roughly $1.46$ million, accounting for approximately $67.3$% of total people in Tokyo at that time. Although this figure is similar to the postal savings, the workers in the manufacturing industry were less likely to rely on the savings banks than the workers in the other industrial sectors. Table A.3 summarizes the number of depositors and the amounts of deposits in savings banks by occupation in Tokyo City (Panel A) and Kyobashi Ward (Panel B). In Kyobashi, the commerce and miscellaneous sectors shared roughly $84.8$% of all the depositors, whereas the manufacturing sector had only $11.5$%. This is unchanged if I use the amounts of deposits to calculate the shares. Tokyo Institute for Municipal Research stated that this tendency captures the fact that savings banks are mainly designed as financial institutions for low-income households (Tokyo Institute for Municipal Research 1925b, p. 81). Thus, skilled factory worker households, like the THBS households, were less likely to use the savings banks. This is consistent with the fact that Tsukishima did not have any savings bank branches, while there were two postal offices. [^88] There were 123 savings banks in Tokyo City in 1922, and 10 of them were in Kyobashi Ward (Tokyo Institute for Municipal Research 1925, p. 76). However, Tsukishima, a representative manufacturing area, did not have savings banks at that time.
@@
Mutual loan association (mujin) is another type of savings institution. The members of this association pay a fixed amount of money several times, and each member withdraws the total amount of payments each time under specific rules. The lotteries or bidding system decided which members could withdraw money each time. This means that the workers could not use this institution to deal with the acute negative income shocks. [^89] This is the most representative mutual loan system called tsūjyō (i.e., common) mujin. There was a different type called sueoki chokin (or tsumitate kai) that people could receive their stakes with interests after a fixed amount had been accumulated for a certain period (Tokyo Legislative Research Association 1915, pp. 3–4). Clearly, both are also not flexible for compensating the income shocks. Instead, it had been used to make a certain amount of money among small business owners and merchants. A report suggests that the workers in the manufacturing sector shared only $10$% of all the mutual loan associations in Tokyo (Tokyo City Office 1935, p. 25). Moreover, the Tokyo Chamber of Commerce stated that the mutual loan association was used for business and rarely for household consumption (Tokyo Chamber of Commerce 1918, p. 121).
@@
Cooperative societies were also available for worker households. Cooperative societies in Tokyo used credit as their payment method: households could buy goods in advance and pay at the end of the month or twice (middle and the end of) a month. The commodity prices were similar to or slightly lower than those in the retail shops (Central Federation of Industrial Associations 1925, pp. 18–19). However, the cooperatives only covered a few citizens. There were 26 cooperative societies in Tokyo prefecture in 1924, and 23 of them were organized by the citizens and workers (Central Federation of Industrial Associations 1925, p. 57–59). Among these 23 cooperatives, two were large societies with roughly 4000 members, and the others were small societies with less than 100 members. The total number of accounts for only a few percentage points in the workers in the city, given that the number of male workers was more than 768 thousand (Table A.1). In fact, most of the THBS households had not used the cooperatives.
@@
An alternative borrowing institution is the credit purchase in retail shops. The THBS households had frequently used credit purchases in various retail shops for rice, fish, vegetables, alcohol, firewood, charcoal, and medicines. The details of the credit purchase are little documented, and the systematic statistics on the credit purchase are also unavailable. This must be because the information on the transactions is personal. Despite this, an available document shows that credit purchases are used in various retail shops (Section 5). The THBS budget data show a consistent tendency for credit purchases to be widely used among factory worker households.
@@
Pawnshops were the popular lending institution among low-income working-class households. Lenders did not need to screen the borrower’s credit, and the borrowers did not worry about incurring heavy debt as the primary articles were inexpensive clothes (Shibuya et al. 1982). In addition, the interest rates were regulated by the Pawnbroker Regulation Act of 1895, and the average redemption rate was substantially high under the lower interest rates. Among $88$ pawnshops in Kyobashi, the average redemption rate was approximately $94$% ($401,222$ cases/$427,265$ cases) in 1920 (Tokyo City Office 1922b, pp. 888–889). The accessibilities were higher than the other lending institutions. The number of pawnshops in Tokyo was $1,334$ in 1918 and $1,261$ in 1919, respectively. The total number of cases was $8,226,883$ and $7,573,406$ each year. These figures are more than three times greater than the entire population in Tokyo City. This means that the accessibility to a pawnshop was relatively high for the workers. As described in the main text, the number of pawnshops in Tsukishima was 13, whereas that of moneylenders was five.
@@
However, the inexpensive article means that pawnshops were used to deal with very short-run necessities only. Among the pawnshops of Kyobashi, for example, the average amount per case was $7.8$ yen, which was $12$% of the monthly earnings of the THBS households. While the pawnshops were popular lending institutions, they could not offer a sufficient amount of temporary income for relatively large idiosyncratic shocks. According to the statistics on the pawnshop users in Tokyo prefecture in 1923, the share of users in the agriculture, commerce, and manufacturing sectors was 1.7%, 19.5%, and 15.6%, respectively. In turn, the most representative users were the day laborers and workers classified as “miscellaneous”, accounting for $63.4$% (Tokyo Institute for Municipal Research 1926, p. 25). In fact, the Tokyo City Office stated that the pawnshops were mainly used among low-income (poor) working-class households (Tokyo City Social Affairs Bureau 1921, p. 9). This indicates the evidence that the primary pawnshop users were not the skilled-worker households like the THBS households.
@@
Another lending institution is the money lenders (kinsen kashitsuke gyō). Generally, business owners used the money lenders. The interest rates in the money lenders were very high compared with those in the pawnshops (Shibuya 2000, pp. 184; 248). Tokyo Institute for Municipal Research was concerned about the increasing number of unscrupulous companies arrested by the National Police Agency (Tokyo Institute for Municipal Research 1925b, pp. 106–108). Given this fact, while there were two money lenders in Tsukishima, they might have been less likely used by the factory worker households.
@@
Credit unions (shinyō kumiai) are another institution available for working-class households. In Tokyo City, there were 215 unions in 1923. According to the survey of the 184 unions among them, however, the number of union members in the manufacturing sector was only $5,662$ (Tokyo Institute for Municipal Research 1925c, p. 20). This accounts for, at the most, only a few percentage points of the male workers in the manufacturing sector in Tokyo measured in the population census of 1920 (Online Appendix Table A.1).
@@
      Number of depositors 76210 550784 166498 619963 1413455
      Amount of deposits 2002124 19112713 6663091 19809591 47587519
@@
      Number of depositors 3809 34665 11664 51481 101619
      Amount of deposits 130781 1218081 450805 1696086 3495753
@@
Notes: This table summarizes the number of depositors and the amount of deposits in the savings banks by occupation. In the document, the miscellaneous category is not clearly defined but named as “Miscellaneous (zatsu gyō)”. These statistics were measured on 31st December 1919. Panel A and B list the statistics for Tokyo City and Kyobashi Word, respectively.
@@ A.3 Development of Tsukishima
Notes: Tsukishima contains Ishikawajima, Tsukudajima, Shin-tsukudajima, and Tsukishima Lots 1–3. This map does not illustrate the northernmost part of Ishikawa jima (see Figure 3 for this part). “Ferries” shows the Tsukishima to Kyobashi Ward sea routes: Tsukuda, Tsukishima, and Kachidoki routes (right to left). “Aioi bridge” links Shin tsukudajima and Fukagawa ward.
@@
Tsukishima (transliterated as “moon island”) is a reclaimed land on a sandbar of the Sumida River. Ishikawajima and Tsukudajima (“jima” means island) at the northern end of Tsukishima are the starting point of its history (Figure A.3 for map). The following description of this section is based on Chūō Ward (1994, pp. 122–124).
@@
Ishikawajima was built in 1626 by a boatman, Hachizaemon Ishikawa, who received a fief from the Shogunate. In 1790, the Ishikawajima Prison was built on Ishikawajima. This was a facility for the reintegration of minor offenders and those without lodgings and was called a Ninsoku Yoseba. After Matthew Perry arrived in Japan in 1853, the Shogunate ordered the Mito Clan to build a shipyard on Ishikawajima. A month later, the groundbreaking ceremony was held for a Western-style wooden sailing ship delivered to the Shogunate in 1856. In 1866, the Ishikawajima Lighthouse was built on the west bank of the Yoseba and served to signal the position of ships entering the Nihonbashi direction. In 1870, the Ninsoku Yoseba became a place of imprisonment (Toba) and was renamed the Ishikawajima Jail Station of the Metropolitan Police Department in 1877. The prison was finally moved to Sugamo in Kitatoshima County in Tokyo prefecture, and the penitentiary on Ishikawajima was closed in 1895.
@@
Adjacent to Ishikawajima is called Tsukudajima, where fishermen from “Tsukuda” in Osaka prefecture settled in 1644 after receiving land from the Shogunate (i.e., the island was named after their hometown, Tsukuda). In 1645, the “Tsukuda ferry” began, and throughout the Edo period, Tsukudajima was known as a fishing island for whitefish for the Shogun’s household. In 1863, the Tsukudajima Battery was built on the south side of Tsukudajima, one of 13 batteries used to defend Edo Bay. This battery later became the starting point for the reclamation of Lot 1 of Tsukishima.
@@
In 1884, the Port of Tokyo Waterway Dredging Project was initiated. This was to secure the passage of ships by sweeping away the sediment in the shipping channel, and the Tsukishima landfill was created using the deposit. The first Tsukishima site was completed in 1891, and the “Tsukishima ferry” began the same year. The second site was created in 1894; the Shin (new) tsukudajima was built in 1896. From the west shore of Shin tsukudajima, the island overlooks the Boso Peninsula in Chiba Prefecture, making it a favorite vacation spot for writers. Cultural figures such as Toson Shimazaki and Kaoru Osanai stayed here from 1907 to the early Taisho period. In 1903, the Aioi Bridge was completed, bringing water and electricity to Tsukishima and improving transportation convenience, which contributed to industry development. [^90] Until then, Suzukiiya had been selling water to supply the entire island using a ferryboat. The “Kachidoki Ferry” began to lead to Lot 2 in 1905. In 1913, Lot 3 was finally completed.
A.4 Housings in Tsukishima
Since the Tsukishima Survey included a housing survey (Table B.1), the housing types in the entire island can be determined. Table A.4 summarizes the results of this survey. The description in this section is based on the Department of Health, Ministry of the Interior (1923a, pp. 69–76).

### v10

@@ 6 Conclusion
This study investigates the consumption-smoothing strategies used by working-class households in prewar Tokyo. I digitized the daily longitudinal budget survey on factory worker households conducted in Tsukishima in 1919. I carefully assessed the potential measurement error due to the miss-assignments of paydays and proposed a way to avoid the assignment errors. The estimated elasticity for total consumption indicates that the factory worker households suffered certain consumption losses when they faced idiosyncratic income shocks.
The estimated elasticity for total expenditure ($0.43$) is in a similar range to those in rural economies in developing countries. According to Ravallion and Chaudhuri (1997), the elasticity for the rural households in Indian villages between 1975 and 1984 was approximately $0.5$. As for the urban households, the estimated elasticity was $0.4$ for Bangkok between 1975 and 1990, which had a similar GDP per capita to 1920 Japan (Townsend 1995). Since my estimate is from a specific occupation, skilled factory workers, we need to be careful when making comparisons with these estimates based on urban samples which include a variety of occupations. Nevertheless, the factory workers in the historical metropolis of Japan circa WWI were vulnerable to idiosyncratic income shocks.
A recent study found an income elasticity of $0.39$ for factory workers’ households in Osaka city circa 1920 (Ogasawara 2024). My estimate from the adjusted monthly panels is slightly larger than this estimate. If taken literally, this implies that the Tsukishima factory workers were slightly more vulnerable than those in Osaka. This indicates that higher levels of urbanization do not necessarily correspond to a greater availability of risk-mitigation institutions. Therefore, inter-city and cross-national comparisons of income elasticities and analysis of the institutional characteristics behind consumption are subjects for future work.
Despite the vulnerability, the results from mechanism analyses show that credit purchases were commonly used among workers and played a role in mitigating adverse idiosyncratic shocks in several consumption subcategories, particularly food. My findings offer a new historical view of the interdependence between consumers and local retailers. As researchers build an understanding of local credit institutions we can comprehend regional consumption behavior and may understand how living standards improved in prewar Japan.
@@ A.1 Development of the Cities
Figure A.1: Development of the Large Cities in Japan
Notes: This figure illustrates the development of large cities in Japan between 1889 and 1920. Figure A.1a shows the number of people in Tokyo, Osaka, Kyoto, and Nagoya cities. Figure A.1b shows the share of people in Tokyo City (diamond) and the cities with more than $50,000$ people (circle) in total population in Japan.
@@
Panel A: The number of people and male workers measured in the Population Census conducted in October 1920 are listed.
Panel B: Occupations are classified based on the industrial classification of the 1920 Population Census. Unemployed people are not included in the calculation.
Panel C: The number of electrical lights measured at the end of 1921 is divided by the number of people in the 1920 Population Census. The number of houses accessible to water supply measured at the end of September 1920 is divided by the number of total households measured in the 1920 Population Census (%).
Panel D: The average daily temperature and precipitation between January 1, 1918, and December 31, 1923, and between February 1, 1920, and April 30, 1920 are listed.
@@ A.2 Financial Institutions in the Cities
Several financial institutions were available for working-class households in the cities circa 1920.
@@
A representative savings institution was the postal savings (yūbin chokin). Postal savings was initially designed as a small savings institution for low-income households, and the share of postal savings to total savings was approximately $10$% in the 1910s (Tanaka 2018, p. 29). However, the number of postal savings accounts per total population in the early 1910s exceeded $20$% (Tanaka 2018, p. 33). James and Suto (2011) indeed show that the saving rate in prewar Japan was higher than that of the US. As shown in Panel A of Table A.2, approximately $1.4$ million people had accounts in Tokyo in 1920. This covers approximately $67$% of the entire population in Tokyo City and the average savings amount per capita was $64.5$ yen. In Kyobashi Ward, the average figure for the postal savings accounts was $44.5$ yen at the end of 1919 (Tokyo City Office 1921, pp. 892–893). This amount was less than the average monthly earnings of the skilled-factory workers (Section 3.2). Unfortunately, the statistics on postal savings by depositors’ occupation are unavailable. [^82] There are some statistics for the whole of Japan. According to the Postal Savings Bureau (1924, p. 72), the share of depositors in the manufacturing sector was $4.8$% (the share based on the deposit amount was $5.5$%). However, this implies that, while a large proportion of the workers might have had accounts, postal savings may not have provided sufficient temporary income.
@@
Savings bank (chochiku ginkō) was another popular savings institution for working-class households in the 1910s (Ito and Saito 2019, pp 79–81). Panel A of Table A.2 shows that the number of people who had savings accounts was approximately $1.46$ million, accounting for approximately $67.3$% of the total people in Tokyo City at that time. Although this figure is similar to the postal savings, the workers in the manufacturing industry were less likely to rely on the savings banks than the workers in the other industrial sectors. Table A.3 summarizes the number of depositors and the amounts of deposits in savings banks by occupation in Tokyo City (Panel A) and Kyobashi Ward (Panel B). In Kyobashi, the commerce and miscellaneous sectors shared approximately $84.8$% of all the depositors, whereas the manufacturing sector had only $11.5$%. This is unchanged if I use the amounts of deposits to calculate the shares. The Tokyo Institute for Municipal Research confirmed that savings banks were designed as financial institutions for low-income households (Tokyo Institute for Municipal Research 1925b, p. 81). Thus, skilled factory worker households—like the THBS households—were less likely to use the savings banks. This is consistent with the fact that Tsukishima did not have any savings bank branches, while there were two postal offices. [^83] There were 123 savings banks in Tokyo City in 1922, and ten of them were in Kyobashi Ward (Tokyo Institute for Municipal Research 1925, p. 76). However, Tsukishima—a representative manufacturing area—did not have savings banks at that time.
@@
Mutual loan association (mujin) was another type of savings institution. The members of this association paid a fixed amount of money several times, and each member withdrew the total amount of payments each time under specific rules. The lotteries or bidding system decided which members could withdraw money each time. This means that the workers could not use this institution to manage acute adverse income shocks. [^84] This is the most representative mutual loan system called tsūjyō (i.e., common) mujin. There was a different type called sueoki chokin (or tsumitate kai) from which people could receive their stakes with interests after a fixed amount had been accumulated for a certain period (Tokyo Legislative Research Association 1915, pp. 3–4). However, these were also not flexible for compensating for income shocks. Instead, these were used by small business owners and merchants to generate money. A report suggests that the workers in the manufacturing sector shared only $10$% of all the mutual loan associations in Tokyo City (Tokyo City Office 1935, p. 25). Moreover, the Tokyo Chamber of Commerce stated that the mutual loan association was used for business and rarely for household consumption (Tokyo Chamber of Commerce 1918, p. 121).
@@
Cooperative societies were also available for worker households. Cooperative societies in Tokyo used credit as their payment method—households could buy goods in advance and pay at the end of the month or twice (in the middle and at the end of) a month. The commodity prices were similar to or slightly lower than those in the retailers (Central Federation of Industrial Associations 1925, pp. 18–19). However, the cooperatives covered a small proportion of citizens. There were $26$ cooperative societies in Tokyo prefecture in 1924, and 23 of them were organized by the citizens and workers (Central Federation of Industrial Associations 1925, p. 57–59). Among these $23$ cooperatives, two were large societies with approximately $4,000$ members, and the others were small societies with less than $100$ members. Given that the number of male workers was more than $768$ thousand (Table A.1), these account for only a few percentage points of the total workers in the city. In fact, the THBS households rarely used the cooperatives.
@@
An alternative borrowing institution was credit purchase in retailers. The details of the credit purchase are little documented, and the systematic statistics on credit purchases are also unavailable. I surmise that this was because the information on the transactions was personal. Despite this, a few available documents show that credit purchases were used in various retailers at that time (Section 5). The THBS households frequently used credit for purchases in various retailers for rice, fish, vegetables, alcohol, firewood, charcoal, and medicines.
@@
Pawnshops were the most popular lending institution among low-income working-class households. Lenders did not need to screen the borrower’s credit, and the borrowers did not worry about incurring heavy debt as the primary articles pawned were inexpensive clothes (Shibuya et al. 1982). In addition, the interest rates were regulated by the Pawnbroker Regulation Act of 1895. Thus, the average redemption rate was substantially high under the lower interest rates. Among $88$ pawnshops in Kyobashi, the average redemption rate was approximately $94$% ($401,222$ cases/$427,265$ cases) in 1920 (Tokyo City Office 1922b, pp. 888–889). The number of pawnshops in Tokyo was $1,334$ and $1,261$ in 1918 and 1919, respectively. The total number of cases was $8,226,883$ and $7,573,406$ each year. These figures were more than three times greater than the entire population in Tokyo City. This means that pawnshops were more accessible to workers than other institutions. As described in the main text, there were $13$ pawnshops in Tsukishima, whereas there were five moneylenders.
@@
However, the inexpensive articles pawned meant that pawnshops were available only for short-run necessities. Among the pawnshops of Kyobashi, for example, the average amount per case was $7.8$ yen, which was approximately $10$% of the monthly earnings of the THBS households. While the pawnshop was a popular lending institution, it could not offer enough temporary income for relatively large idiosyncratic shocks. According to the statistics on the pawnshop users in Tokyo prefecture in 1923, the share of users in the agriculture, commerce, and manufacturing sectors was $1.7$%, $19.5$%, and $15.6$%, respectively. In turn, the most representative users were the day laborers and workers classified as “miscellaneous,” accounting for $63.4$% (Tokyo Institute for Municipal Research 1926, p. 25). In fact, the Tokyo City Office stated that the pawnshops were mainly used among low-income (poor) working-class households (Tokyo City Social Affairs Bureau 1921, p. 9). This indicates the evidence that the primary pawnshop users were not the skilled-worker households like the THBS households.
@@
Another lending institution was the money lenders (kinsen kashitsuke gyō). Generally, business owners used money lenders. The interest rates of the money lenders were very high compared with those in the pawnshops (Shibuya 2000, pp. 184; 248). Tokyo Institute for Municipal Research was indeed concerned about the increasing number of unscrupulous companies arrested by the National Police Agency (Tokyo Institute for Municipal Research 1925b, pp. 106–108). Given this fact, while there were two money lenders in Tsukishima, they were less likely to be used by the factory worker households.
@@
Credit union (shinyō kumiai) was another institution available for working-class households. In Tokyo City, there were $215$ unions in 1923. According to the survey of the $184$ unions among them, however, the number of union members in the manufacturing sector was only $5,662$ (Tokyo Institute for Municipal Research 1925c, p. 20). This accounts for only a few percentage points of the male workers in the manufacturing sector in Tokyo measured in the population census of 1920 (Online Appendix Table A.1).
@@
      Number of depositors 76,210 550,784 166,498 619,963 1,413,455
      Amount of deposits 2,002,124 19,112,713 6,663,091 19,809,591 47,587,519
@@
      Number of depositors 3,809 34,665 11,664 51,481 101,619
      Amount of deposits 130,781 1218,081 450,805 1,696,086 3,495,753
@@
Notes: This table summarizes the number of depositors and the total deposits in the savings banks by occupation. In the document, the miscellaneous category is not clearly defined but named as “Miscellaneous (zatsu gyō).” These statistics were measured on December 31, 1919. Panel A and B list the statistics for Tokyo City and Kyobashi Word, respectively.
@@ A.3 Development of Tsukishima
Notes: Tsukishima contains Ishikawajima, Tsukudajima, Shin-tsukudajima, and Tsukishima Lots 1–3. This map does not illustrate the northernmost part of Ishikawa jima (see Figure 3 for this part). “Ferries” shows the Tsukishima to the mainland of Kyobashi Ward sea routes called Tsukuda, Tsukishima, and Kachidoki routes (right to left). “Aioi bridge” links Shin tsukudajima and Fukagawa ward.
@@
Tsukishima (transliterated as “moon island”) is a reclaimed land on a sandbar of the Sumida River. Ishikawajima and Tsukudajima islands at the northern end of Tsukishima (Figure A.3) are the starting point of its history. The following description in this section is based on Chūō Ward (1994, pp. 122–124).
@@
Ishikawajima was built in 1626 by a boatman, Hachizaemon Ishikawa, who received a fief from the Shogunate. In 1790, the Ishikawajima Prison (ninsoku yoseba) was built on Ishikawajima. This was a facility for the reintegration of minor offenders and those without lodgings. After Matthew Perry arrived in Japan in 1853, the Shogunate ordered the Mito Clan to build a shipyard on Ishikawajima. A month later, a groundbreaking ceremony was held for a Western-style wooden sailing ship delivered to the Shogunate in 1856. In 1866, the Ishikawajima Beacon was built on the west bank of the Ishikawajima Prison and served to signal the position of ships entering the Nihonbashi direction. In 1870, the Ishikawajima Prison became a place of imprisonment (toba) and was renamed the Ishikawajima Jail Station of the Metropolitan Police Department in 1877. The prison was finally moved to Sugamo in Kitatoshima County in Tokyo prefecture, and the penitentiary on Ishikawajima was closed in 1895.
@@
The area adjacent to Ishikawajima is called Tsukudajima, where fishermen from “Tsukuda” in Osaka prefecture settled in 1644 after receiving land from the Shogunate. In 1645, the “Tsukuda ferry” began, and throughout the Edo period, Tsukudajima was known as a fishing island for whitefish for the Shogun’s household. In 1863, the Tsukudajima Battery was built on the south side of Tsukudajima—one of $13$ batteries used to defend Edo Bay. This battery later became the starting point for the reclamation of Lot 1 of Tsukishima.
@@
In 1884, the Port of Tokyo Waterway Dredging Project was initiated. This was to secure the passage of ships by sweeping away the sediment in the shipping channel, and the Tsukishima landfill was created using the deposit. The first Tsukishima site was completed in 1891, and the “Tsukishima ferry” began the same year. The second site was created in 1894; the Shin Tsukudajima (namely, new Tsukudajima) was built in 1896. From the west shore of Shin Tsukudajima, the island overlooks the Boso Peninsula in Chiba Prefecture, making it a favorite vacation spot for writers. Cultural figures such as Toson Shimazaki and Kaoru Osanai stayed here from 1907 to the early Taisho period. In 1903, the Aioi Bridge was completed, bringing water and electricity to Tsukishima and improving transportation convenience, which contributed to development of industry. [^85] Until then, a retailer named Suzukiiya had been selling water to supply the entire island using a ferryboat. The “Kachidoki Ferry” began to deliver to Lot 2 in 1905. Lot 3 was finally completed in 1913.
A.4 Housing in Tsukishima